# Recursive Quantum Relaxation for Combinatorial Optimization Problems

### v2

@@ 4 Numerical Experiments
The second dataset is the benchmark graph dataset Gset [15] (http://www.stanford.edu/~yyye/yyye/Gset/). Gset includes random, toroidal, and almost planar graphs with $|V|=800$ to $20000$ indexed by G1 to G81, which are generated by rudy. Only G1, G6, G14, and G18 are used for evaluation, all of which have 800 nodes. G1 and G6 are random graphs, G11 is a toroidal graph, and G14 and G18 are the combined graphs of two planar graphs. Graphs G1 and G6, and graphs G14 and G18 have the same edge structure except for the weights, where G1 and G14 have $+1$ edge weights whereas G6 and G18 have edge weights of ±$\pm 1$ assigned at random. The edge weights of G11 are ±$\pm 1$ assigned at random. We call this graph dataset Gset800.
@@
We compared our model with GW [7], CirCut [16], QRAO [11], and RQAOA [3]. GW is a well-known classical approximation algorithm based on semidefinite programming. CirCut is the Fortran program name of the rank-two relaxation algorithm that is a classical heuristic algorithm and empirically outperforms other classical heuristics on various graph instances [20]. The details of GW and CirCut are described in Sec. 5.1. The RQAOA is a QAOA-based algorithm whose details are explained in Sec. 2.2. QRAO is a quantum relaxation-based optimization method whose details are explained in Sec. 2.3.
@@
Using the Gset800 dataset, the procedure was run 10 times on the same graph instance and the best cut of the trials was recorded. RQRAO and all baseline methods were used. We call this experiment ExpGset.
@@ 4.1 Results and Discussion
Figure 5 summarizes the results of ExpVer. If one can obtain the eigenstate with the maximum eigenvalue, $(1,1)$-QRAC is the best choice because the cut weight with the highest probability for $(1,1)$-QRAC is always the MAX-CUT solution (Fig. 5(a)). By contrast, in the case of $(2,1)$ and $(3,1)$-QRACs, the solution with the highest probability is sometimes not the MAX-CUT solution. This occurs when low cut weight solutions are embedded near the MAX-CUT solution, and higher but not the maximum cut weight solutions are aggregated somewhere else in Hilbert space. In such a case, the state with the maximum eigenvalue is placed in its aggregation, which could be far from the MAX-CUT solution. This never occurs in the case of $(1,1)$-QRAC, because all embeddings are orthogonal to each other. There are drawbacks due to these accidental cases. However, $(2,1)$ and $(3,1)$-QRACs still have relatively higher cut weights. Therefore, it makes sense to choose a sample with the highest probability as a solution to obtain a higher cut weight solution, which is what the RQAOA, QRAO, and RQRAO attempt to implement according to Theorem 1, if the state corresponding to the maximum eigenvalue can be obtained. However, in practice, the state obtained through optimization is not the eigenstate with the maximum eigenvalue. In this case, the cut weight with the highest probability for both $(2,1)$- and $(3,1)$-QRACs is much better than that for $(1,1)$-QRAC (Fig. 5(b)). This suggests that $(2,1)$- or $(3,1)$-QRAC is a better choice for obtaining a higher cut weight solution when the state is obtained through optimization. One reason for this difference comes from the difficulty of optimization. Optimization using the $(2,1)$- and $(3,1)$-QRAC Hamiltonians would be easier than using the $(1,1)$-QRAC because the numbers of qubits of $(2,1)$- and $(3,1)$-QRAC Hamiltonians are less than that of $(1,1)$-QRAC. The other reason is the difference in the landscape of the objective functions. However, the details have not yet been clarified and are left for future work.
Originally, the MAX-CUT problem in an Ising-type formulation was to find the eigenstate with the maximum eigenvalue, which is known to be QMA-hard [12]. Remember that the variable with the highest probability is embedded in the magic state with the highest fidelity to the current quantum state. This suggests that the MAX-CUT problem has now been changed into a problem of finding the magic state with the highest fidelity to the optimized quantum state. Of course, finding the magic state with the largest fidelity is also intractable, and the largest fidelity sample is not always the MAX-CUT solution, as seen in Fig. 5(a). However, in the practical case, we can expect that a better solution can be obtained using $(2,1)$- and $(3,1)$-QRACs rather than using $(1,1)$-QRAC, as shown in Fig. 5(b). In other words, the MAX-CUT problem could be successfully relaxed by the QRAC Hamiltonian.
Gset (ExpGset)
@@
As expected in Sec. 3.2.2, the cut weight of RQRAO with $N=1$ is very similar to that of QRAO because both assume all random variables are independent. The cut weight performance of RQRAO increases when the number of ensembles $N$ is increased, which confirms the effectiveness of using an ensemble, which is heuristically introduced in our method.
@@
Figure 6 summarizes the results on $3$-regular graphs. Although the relative cuts of QRAO and the RQAOA are lower than 1, that of RQRAO exceeds 1 when the number of nodes is larger than 200 and is larger than that of CirCut (Fig. 6). In addition, although the absolute runtime of RQRAO is longer than that of CirCut, the order of empirical time complexity of these methods is close (Fig. 6). To quantitatively compare the empirical time complexity of the methods, the runtime curves were fitted using
${({\rm Wall\ time\ [sec]})=a({\rm Number\ of\ nodes})^{b}}$
using only data with a large number of nodes. The fitted values of $b$ are $1.34$ for CirCut, $2.78$ for the RQAOA, $1.73$ for QRAO, $1.44$ for RQRAO, and $2.44$ for GW. That is, the empirical time complexity of the proposed method is much lower than that of GW and the RQAOA. Moreover, it is comparable to that of CirCut. Because most of the computational time is spent on optimizing the variational states, the computational time of the proposed method can be reduced by improving the optimization algorithm. For example, the runtime cost of RQRAO can be reduced by parallel computation in the ensemble part. Further reductions in the computational costs of RQRAO are future work.
@@
Only the results for varying the number of embeddings per qubit $m$ are described here. Other hyperparameter sensitivities are summarized in Supplemental Information F. In Fig. 7, we can see the clear advantage of using $m>1$ over $m=1$, which is that only the cut weight in the cases of $m>1$ overcomes that of GW. When $m=1$, the optimized state may be stacked into a bad local minimum, as shown in Fig. 5(b), and this is the reason for the lower cut weight. Almost no runtime dependency on $m$ is seen in Fig. 7, which means there is no disadvantage to using $m>1$ as long as the given graph is sufficiently sparse and the MPS ansatz is adopted.
5 Related Work
5.1 Classical Algorithm for the MAX-CUT problem
$\max_{{\bf X}}\texttt{CW}_{\rm X}({\bf X}),$ (44)
where
$\texttt{CW}_{\rm X}({\bf X}):={\rm tr}({\bf w}{\bf F}({\bf X})),$ (45)
The best known theoretically guaranteed generic algorithm for the MAX-CUT problem is the GW [7] as stated in Sec. 1. In the GW algorithm, the constraint that ${\rm rank}{(\bf X})=1$ is removed. This transforms the MAX-CUT problem to a semidefinite program (SDP), which can be solved in polynomial time for the number of nodes. After solving the SDP, the solution ${\bf v}\in\{\pm 1\}^{|V|}$ needs to be extracted from the obtained ${\bf X}$. When ${\bf X}$ is written in the form ${\bf X}={\bf vv}^{\mathsf{T}}$, the solution is ${\bf v}$. If not, as is generally the case here, the obtained ${\bf X}$ is decomposed as ${\bf X}={\bf L}{\bf L}^{\mathsf{T}}$ using, for example, Cholesky decomposition, where ${\bf L}=\begin{bmatrix}{\bf l}_{1}&\dots&{\bf l}_{|V|}\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{|V|\times|V|}$ is a lower triangular matrix and ${\bf l}_{i}\in{\mathbb{R}}^{|V|}$ is a unit vector. Then, the solution $\{v_{i}\}$ is determined by the random projection of ${\bf l}_{i}$ as $v_{i}=1$ if ${\bf l}_{i}^{\mathsf{T}}{\bf r}\geq 0$ and $v_{i}=-1$ otherwise. Here, ${\bf r}\in{\mathbb{R}}^{|V|}$ is a random vector uniformly distributed on the $|V|$-dimensional unit sphere. This can be viewed as dividing the set $\{{\bf l}_{i}\}^{|V|}_{i=1}$ into two sets by a hyperplane with ${\bf r}$ as the normal vector. When the hyperplane cutting is performed uniformly at random, the lower bound of the expected cut weight is $0.878\,\texttt{CW}({\bf b}^{*})$ [7]. This bound is explained as follows. According to [7, Theorem 3.1], the MAX-CUT problem can be further reformulated as
$\{{\bf l}^{\rm opt}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (46)
where
$\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}):=\frac{{\rm arccos}({\bf l}_{j}^{\mathsf{T}}{\bf l}_{k})}{\pi}.$ (47)
In the GW algorithm, the objective is rewritten as
$\{{\bf l}^{*}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (48)
where
$\leq\frac{1}{0.878}\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}).$ (49)
The last inequality of Eq. (49) comes from [7, Lemma 3.4, Lemma 3.5]. 222 $\min_{x\in[-1,1]}\frac{{\rm arccos}(x)}{\pi}\big{(}\frac{1-x}{2}\big{)}^{-1}\approx 0.878$. The right-hand side of Eq. (48) is an SDP, and hence it can be solved in polynomial time. The resulting ${\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})}[\texttt{CW}({\bf b})]$, which is the expected cut weight of random hyperplane cutting, is guaranteed to be greater than $0.878\,{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{\rm opt}_{i}\})}[\texttt{CW}({\bf b})]=0.878\max_{{\bf b}}\texttt{CW}({\bf b})$ because
${\geq 0.878\,\mathcal{P}^{\rm GW}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\})}$
$=0.878\,\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\}).$ (50)
Equation (48) is very similar to Eq. (18), but the rounding process of the GW is not the same as that of Eq. (19). Instead of carrying out ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}^{*}_{i}\})$ or ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$, a number of ${\bf b}$ obeying $\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$ are sampled by random hyperplane cutting and then, the best $\texttt{CW}({\bf b})$ is adopted as the candidate solution.
$\{\theta^{*}_{j}\}:=\mathop{\rm argmax}_{\{\theta_{j}\}}\mathbb{E}_{\mathcal{P}^{\rm R2}({\bf b};\{\theta_{j}\})}[\texttt{CW}({\bf b})],$ (51)
where
$\mathcal{P}^{\rm R2}(b_{j}\neq b_{k};\{\theta_{j}\})=\sin^{2}\Big{(}\frac{\theta_{j}-\theta_{k}}{2}\Big{)}.$ (52)
${\bf b}^{\rm R2}=\texttt{CW}({\bf b}^{*}(\alpha^{*})),$ (53)
where
$b^{*}_{j}(\alpha)=\left\{\begin{array}[]{lll}0&{\rm for}&\cos(\theta^{*}_{j}+\alpha)\geq 0\\[5.69054pt] 1&{\rm for}&\cos(\theta^{*}_{j}+\alpha)<0\\ \end{array}\right.$ (54)
and
$\alpha^{*}=\mathop{{\rm argmax}}_{\alpha}\texttt{CW}({\bf b}^{*}(\alpha)).$ (55)
5.2 Quantum Algorithms for the MAX-CUT problem
5.3 QRACs
The concept that encodes $m$ bits into $n$ (qu)bits was originally proposed by Wiesner as conjugate coding [45] and later rediscovered by Ambainis et al. [1] as (quantum) random access codes ((Q)RACs). $(m,n,p)$ random access encoding is defined as the function that maps $m$ classical bits into $n$ (qu)bits with a decoding probability at least $p$ [1, Definition 1.1]. The minimum number of $n$ with $p>1/2$ is bounded as $(1-\eta(p))m$ for both classical [1, Theorem 2.1] and quantum [46, Theorem 2.3] settings, where $\eta(p)$ is the binary entropy function. In the case of $(2,1,p)$- and $(3,1,p)$-QRACs, there exists $p=\frac{1}{2}+\frac{1}{2\sqrt{2}}$ [1, Lemma 3.1] and $p=\frac{1}{2}+\frac{1}{2\sqrt{3}}$ [1, attributed to Chuang], respectively, which are consistent with our results (Eq. (30)) in the case of $\mathcal{E}_{j}\in\{\pm 1\}$. It was proved that $(2^{n},n,>1/2)$ classical random access encoding [47, Theorem 5.5] and $(2^{2n},n,>1/2)$ quantum random access encoding [48, Theorem 2] do not exist. From these observations, $(3,1)$-QRAC shows the maximum number of classical bits that can be embedded in one qubit with a decoding probability that is larger than $1/2$. This is why we only consider $m\in\{1,2,3\}$ in this study. In addition to $(m,1,p)$-QRAC, $(m,2,p)$-QRAC is particularly of interest and well studied in terms of the theoretical bound [48, 49, 50, 51] and concrete construction scheme [52]. For arbitrary $(m,n,p)$-QRAC, the upper bound of decoding probability $p$ is improved by considering shared randomness [53], shared entanglement [54], and $d$-level system [55, 49]. The information that can be decoded is not limited to the $m$ bits that are encoded, but also extends to functions that use $k$-out-of-$m$ bits [56, 57]. Using this extension, QRACs have also been applied in machine learning to efficiently encode discrete features [58].
@@ Acknowledgement
Appendix A Classical Shadow with Random Magic Measurements
@@
where $\hat{b}^{[q]}$ is the measurement outcome of the $q$th qubit. Equation (S.3) is the same as that of random Pauli measurements except for the unitary ensemble $\mathcal{U}_{\rm ms}$. In the case of random magic measurements, the unitary ensemble is $\mathcal{U}_{\rm ms}=\{U_{i}\}^{4}_{i=1}$, where $\mu_{i}^{+}:=U_{i}|0\rangle\langle 0|U_{i}^{\dagger}$ and $\mu_{i}^{-}:=U_{i}|1\rangle\langle 1|U_{i}^{\dagger}$. Note that $U_{1}$ can be expressed as $U_{1}^{\dagger}=e^{{\rm i}tX}e^{{\rm i}sZ}$ where $s=\pi/8$ and $t={\rm arccos}(\sqrt{(1+1/\sqrt{3})/2})$, and $U_{2}^{\dagger}=U^{\dagger}_{1}X$, $U^{\dagger}_{3}=U^{\dagger}_{1}Y$, $U^{\dagger}_{4}=U^{\dagger}_{1}Z$ [11].

### v3

@@ 4 Numerical Experiments
The second dataset is the benchmark graph dataset Gset [15] (http://www.stanford.edu/~yyye/yyye/Gset/). Gset includes random, toroidal, and almost planar graphs with $|V|=800$ to $20000$ indexed by G1 to G81, which are generated by rudy. Only G1, G6, G11, G14, and G18 are used for evaluation, all of which have 800 nodes. G1 and G6 are random graphs, G11 is a toroidal graph, and G14 and G18 are the combined graphs of two planar graphs. Graphs G1 and G6, and graphs G14 and G18 have the same edge structure except for the weights, where G1 and G14 have $+1$ edge weights whereas G6 and G18 have edge weights of ±$\pm 1$ assigned at random. The edge weights of G11 are ±$\pm 1$ assigned at random. We call this graph dataset Gset800.
@@
We compared our model with GW [7], CirCut [16], QRAO [11], and RQAOA [3]. GW is a well-known classical approximation algorithm based on semidefinite programming. CirCut is the Fortran program name of the rank-two relaxation algorithm; it is a classical heuristic algorithm that empirically outperforms other classical heuristics on various graph instances [22]. The details of GW and CirCut are described in Supplementary Information H.1. The RQAOA is a QAOA-based algorithm whose details are explained in Sec. 2.2. QRAO is a quantum relaxation-based optimization method whose details are explained in Sec. 2.3.
@@
Using the Gset800 dataset, the procedure was run 10 times on the same graph instance and the best cut weight of the trials was recorded. RQRAO and all baseline methods were used. We call this experiment ExpGset.
@@ 4.1 Results and Discussion
Figure 6 summarizes the results of ExpVer. When the eigenstate with the maximum eigenvalue is obtained, the $(1,1)$-QRAC is consistently the best choice as it provides the optimal MAX-CUT solution with the highest probability (Fig. 6(a)). On the other hand, for $(2,1)$ and $(3,1)$-QRACs, particularly when higher cut weight solutions cluster in a different location within the Hilbert space from the MAX-CUT solution, the solution with the highest probability may not correspond to the optimal MAX-CUT solution. Therefore, even if the ground state is obtained, there is a disadvantage that the nearest solution may not be the optimal MAX-CUT solution.
Nevertheless, when the quantum state is obtained through optimization (Fig. 6(b)), $(2,1)$ and $(3,1)$-QRACs achieve higher cut weights than when using $(1,1)$-QRAC. The reason why higher cut weights can be obtained by setting $m\geq 2$ can be explained as follows. As shown in Fig. 1, as $m$ increases, more solutions are packed into fewer qubits, resulting in an average over a greater number of solutions being observed. In other words, as $m$ increases, we can only observe the average of a larger number of solutions, leading to a blurred expectation value. Since the optimization process attempts to reach the optimal state from such a blurred function value that only reflects the average, increasing $m$ does not necessarily bring the optimized state closer to the nearest state embedding the optimal solution. However, this blurring could also potentially make the function value smoother, thus simplifying the optimization process. Such optimization-related advantages occur for $m=2$ and $m=3$, and it is likely that $m=2$ resulted in the highest cut weight because it strikes a balance between the aforementioned advantages and disadvantages.
In this study, we have optimized a classically trackable MPS ansatz using classical optimization methods. However, recent developments, such as quantum state preparation using Lindbladians [26], suggest that pure quantum algorithms are promising. In our method, more than 90% of the computation time is spent on parameter optimization, so there could be benefits in terms of overall algorithm speedup by replacing the optimization process with a pure quantum algorithm. However, since the discussion on optimization is beyond the scope of this study, these considerations will be addressed in future work. In any case, the MAX-CUT problem, traditionally a problem of finding the eigenstate with the maximum eigenvalue, has effectively been shifted to finding the magic state embedding with the highest fidelity to the optimized quantum state. While challenging, this approach may yield better solutions, particularly with $(2,1)$ and $(3,1)$-QRACs. In the following, we will explore the advantages of $m\geq 2$ through additional experiments.
Gset (ExpGset)
@@
As expected in Sec. 3.2.2, the cut weight of RQRAO with $N=1$ is very similar to that of QRAO because both assume all random variables are independent. The cut weight performance of RQRAO increases when the number of ensembles $N$ is increased, which confirms the effectiveness of using an ensemble.
@@
Figure 7 summarizes the results on $3$-regular graphs. Although the relative cut weights of QRAO and the RQAOA are lower than 1, that of RQRAO exceeds 1 when the number of nodes is larger than 200 and is larger than that of CirCut (Fig. 7). In addition, although the absolute runtime of RQRAO is longer than that of CirCut, the order of empirical time complexity of these methods is close (Fig. 7). To quantitatively compare the empirical time complexity of the methods, the runtime curves were fitted using
${({\rm Wall\ time\ [sec]})=\alpha({\rm Number\ of\ nodes})^{\beta}}$
using only data with a large number of nodes. The fitted values of $\beta$ are $1.34$ for CirCut, $2.71$ for the RQAOA, $1.73$ for QRAO, $1.44$ for RQRAO, and $2.44$ for GW. That is, the empirical time complexity of the proposed method is much lower than that of GW and the RQAOA, and comparable to that of CirCut. Because most of the computational time is spent on optimizing the variational states, the computational time of the proposed method can be reduced by improving the optimization algorithm. For example, the runtime cost of RQRAO can be reduced by parallel computation in the ensemble part. Further reductions in the computational costs of RQRAO are future work. In Supplementary Information G, we also compared the methods using the 2D toric graphs with one extra node connected to all other nodes, as used in [3] to demonstrate the power of RQAOA. The results reaffirm what we have observed when comparing the methods using 3-regular graphs.
@@
Only the results for varying the number of embeddings per qubit $m$ are described here. Other hyperparameter sensitivities such as the number of ensemble $N$, the scale factor $S$, and the bond dimension $\chi$, are summarized in Supplemental Information F. In Fig. 8, we can see the clear advantage of using $m>1$ over $m=1$; only the cut weight in the cases of $m>1$ overcomes that of GW. When $m=1$, the optimized state may be stacked in bad local optima, as shown in Fig. 6(b), and this is the reason for the lower cut weight. Almost no runtime dependency on $m$ is seen in Fig. 8, which means there is no disadvantage to use $m>1$ as long as the given graph is sufficiently sparse to allow applying $(m,1)$-QRAC Hamiltonian.
@@ Acknowledgement
Appendix A Random Magic Measurements and the Relationship with Classical Shadows
@@
where $\hat{b}^{[q]}$ is the measurement outcome of the $q$th qubit. Equation (S.3) is the same as that of random Pauli measurements except for the unitary ensemble $\mathcal{U}_{\rm ms}$. In the case of random magic measurements, the unitary ensemble is $\mathcal{U}_{\rm ms}=\{U_{i}\}^{4}_{i=1}$, where $\mu_{i}^{+}:=U_{i}|0\rangle\langle 0|U_{i}^{\dagger}$ and $\mu_{i}^{-}:=U_{i}|1\rangle\langle 1|U_{i}^{\dagger}$. Note that $U_{1}$ can be expressed as $U_{1}^{\dagger}=e^{-{\rm i}tX}e^{-{\rm i}sZ}$ where $s=\pi/8$ and $t={\rm arccos}(\sqrt{(1+1/\sqrt{3})/2})$, and $U_{2}^{\dagger}=U^{\dagger}_{1}X$, $U^{\dagger}_{3}=U^{\dagger}_{1}Y$, $U^{\dagger}_{4}=U^{\dagger}_{1}Z$ [11].
@@ Appendix E Implementation Details for GW, CirCut, the RQAOA, and QRAO
Figure S.4: (a) Example of 2D toric graph with one external nodes of size $3\times 3$ (# girds / dimension $=3$); (b) Summary of results for 2D-toric graphs.
@@ Appendix F Additional Sensitivity Analysis Results (ExpHyp)
Appendix G Results for 2D Toric Graphs
While the specific characteristics of instances where RQAOA performs particularly well are not fully understood, in [3] RQAOA was compared against GW using 2D toric graphs with one extra node connected with all other nodes (Fig. S.4), referred to as $G_{\rm toric}$. To verify that our proposed model also surpasses RQAOA on $G_{\rm toric}$, we conducted numerical simulations for $G_{\rm toric}$ under the same settings as ExpScale. In Fig. S.4, it is evident that RQAOA outperforms GW in terms of cut weight when the number of grids per dimension exceeds 15, consistent with the findings of [3]. However, both CirCut and our proposed RQRAO surpass RQAOA. Note that the trend in time complexity is nearly identical to the results observed for 3-regular graphs, as shown in Fig. 7.
Appendix H Related Work
H.1 Classical Algorithm for the MAX-CUT problem
$\max_{{\bf X}}\texttt{CW}_{\rm X}({\bf X}),$ (S.31)
where
$\texttt{CW}_{\rm X}({\bf X}):={\rm tr}({\bf w}{\bf F}({\bf X})),$ (S.32)
The best known theoretically guaranteed generic algorithm for the MAX-CUT problem is the GW [7] as stated in Sec. 1. In the GW algorithm, the constraint ${\rm rank}{(\bf X})=1$ is removed. This transforms the MAX-CUT problem to a semidefinite program (SDP), which can be solved in polynomial time for the number of nodes. After solving the SDP, the solution ${\bf v}\in\{\pm 1\}^{|V|}$ needs to be extracted from the obtained ${\bf X}$. When ${\bf X}$ is written in the form ${\bf X}={\bf vv}^{\mathsf{T}}$, the solution is ${\bf v}$. If not, as is generally the case here, the obtained ${\bf X}$ is decomposed as ${\bf X}={\bf L}{\bf L}^{\mathsf{T}}$ using, for example, Cholesky decomposition, where ${\bf L}=\begin{bmatrix}{\bf l}_{1}&\dots&{\bf l}_{|V|}\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{|V|\times|V|}$ is a lower triangular matrix and ${\bf l}_{i}\in{\mathbb{R}}^{|V|}$ is a unit vector. Then, the solution $\{v_{i}\}$ is determined by the random projection of ${\bf l}_{i}$ as $v_{i}=1$ if ${\bf l}_{i}^{\mathsf{T}}{\bf r}\geq 0$ and $v_{i}=-1$ otherwise. Here, ${\bf r}\in{\mathbb{R}}^{|V|}$ is a random vector uniformly distributed on the $|V|$-dimensional unit sphere. This can be viewed as dividing the set $\{{\bf l}_{i}\}^{|V|}_{i=1}$ into two sets by a hyperplane with ${\bf r}$ as the normal vector. When the hyperplane cutting is performed uniformly at random, the lower bound of the expected cut weight is $0.878\,\texttt{CW}({\bf b}^{*})$ [7]. This bound is explained as follows. According to [7, Theorem 3.1], the MAX-CUT problem can be further reformulated as
$\{{\bf l}^{\rm opt}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (S.33)
where
$\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}):=\frac{{\rm arccos}({\bf l}_{j}^{\mathsf{T}}{\bf l}_{k})}{\pi}.$ (S.34)
In the GW algorithm, the objective is rewritten as
$\{{\bf l}^{*}_{i}\}:=\mathop{\rm argmax}_{\{{\bf l}_{i}\}}{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}_{i}\})}[\texttt{CW}({\bf b})]$ (S.35)
where
$\leq\frac{1}{0.878}\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}_{i}\}).$ (S.36)
The last inequality of Eq. (S.36) comes from [7, Lemma 3.4, Lemma 3.5]. 222 $\min_{x\in[-1,1]}\frac{{\rm arccos}(x)}{\pi}\big{(}\frac{1-x}{2}\big{)}^{-1}\approx 0.878$. The right-hand side of Eq. (S.35) is an SDP, and hence it can be solved in polynomial time. The resulting ${\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})}[\texttt{CW}({\bf b})]$, which is the expected cut weight of random hyperplane cutting, is guaranteed to be greater than $0.878\,{\mathbb{E}}_{{\bf b}\sim\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{\rm opt}_{i}\})}[\texttt{CW}({\bf b})]=0.878\max_{{\bf b}}\texttt{CW}({\bf b})$ because
${\geq 0.878\,\mathcal{P}^{\rm GW}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\})}$
$=0.878\,\mathcal{P}^{\rm MC}(b_{j}\neq b_{k};\{{\bf l}^{\rm opt}_{i}\}).$ (S.37)
Equation (S.35) is very similar to Eq. (19), but the rounding process of the GW is not the same as that of Eq. (20). Instead of carrying out ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm GW}({\bf b};\{{\bf l}^{*}_{i}\})$ or ${\bf b}^{\rm R}=\mathop{\rm argmax}_{{\bf b}}\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$, a number of ${\bf b}$ obeying $\mathcal{P}^{\rm MC}({\bf b};\{{\bf l}^{*}_{i}\})$ are sampled by random hyperplane cutting and then, the best $\texttt{CW}({\bf b})$ is adopted as the candidate solution.
$\{\theta^{*}_{j}\}:=\mathop{\rm argmax}_{\{\theta_{j}\}}\mathbb{E}_{\mathcal{P}^{\rm R2}({\bf b};\{\theta_{j}\})}[\texttt{CW}({\bf b})],$ (S.38)
where
$\mathcal{P}^{\rm R2}(b_{j}\neq b_{k};\{\theta_{j}\})=\sin^{2}\Big{(}\frac{\theta_{j}-\theta_{k}}{2}\Big{)}.$ (S.39)
${\bf b}^{\rm R2}=\texttt{CW}({\bf b}^{*}(\alpha^{*})),$ (S.40)
where
$b^{*}_{j}(\alpha)=\left\{\begin{array}[]{lll}0&{\rm for}&\cos(\theta^{*}_{j}+\alpha)\geq 0\\[5.69054pt] 1&{\rm for}&\cos(\theta^{*}_{j}+\alpha)<0\\ \end{array}\right.$ (S.41)
and
$\alpha^{*}=\mathop{{\rm argmax}}_{\alpha}\texttt{CW}({\bf b}^{*}(\alpha)).$ (S.42)
H.2 Quantum Algorithms for the MAX-CUT problem
H.3 QRACs
The concept that encodes $m$ bits into $n$ (qu)bits was originally proposed by Wiesner as conjugate coding [51] and later rediscovered by Ambainis et al. [1] as (quantum) random access codes ((Q)RACs). $(m,n,p)$ random access encoding is defined as the function that maps $m$ classical bits into $n$ (qu)bits with a decoding probability at least $p$ [1, Definition 1.1]. The minimum number of $n$ with $p>1/2$ is bounded as $(1-\eta(p))m$ for both classical [1, Theorem 2.1] and quantum [52, Theorem 2.3] settings, where $\eta(p)$ is the binary entropy function. In the case of $(2,1,p)$- and $(3,1,p)$-QRACs, there exists $p=\frac{1}{2}+\frac{1}{2\sqrt{2}}$ [1, Lemma 3.1] and $p=\frac{1}{2}+\frac{1}{2\sqrt{3}}$ [1, attributed to Chuang], respectively, which are consistent with our results (Eq. (28)) in the case of $\mathcal{E}_{j}\in\{\pm 1\}$. It was proved that $(2^{n},n,>1/2)$ classical random access encoding [53, Theorem 5.5] and $(2^{2n},n,>1/2)$ quantum random access encoding [54, Theorem 2] do not exist. From these observations, $(3,1)$-QRAC shows the maximum number of classical bits that can be embedded in one qubit with a decoding probability that is larger than $1/2$. This is why we only consider $m\in\{1,2,3\}$ in this study. In addition to $(m,1,p)$-QRAC, $(m,2,p)$-QRAC is particularly of interest and well studied in terms of the theoretical bound [54, 55, 56, 57] and concrete construction scheme [58]. For arbitrary $(m,n,p)$-QRAC, the upper bound of decoding probability $p$ is improved by considering shared randomness [59], shared entanglement [60], and $d$-level system [61, 55]. The information that can be decoded is not limited to the $m$ bits that are encoded, but also extends to functions that use $k$-out-of-$m$ bits [62, 63]. Using this extension, QRACs have also been applied in machine learning to efficiently encode discrete features [64].